### v1

# Assortment Optimization for Patient-Provider Matching

Abstract
Rising provider turnover forces healthcare administrators to frequently rematch patients to available providers, which can be cumbersome and labor-intensive. To reduce the burden of rematching, we study algorithms for matching patients and providers through assortment optimization. We develop a patient-provider matching model in which we simultaneously offer each patient a menu of providers, and patients subsequently respond and select providers. By offering assortments upfront, administrators can balance logistical ease and patient autonomy. We study policies for assortment optimization and characterize their performance under different problem settings. We demonstrate that the selection of assortment policy is highly dependent on problem specifics and, in particular, on a patient’s willingness to match and the ratio between patients and providers . On real-world data, we show that our best policy can improve match quality by 13% over a greedy solution by tailoring assortment sizes based on patient characteristics. We conclude with recommendations for running a real-world patient-provider matching system inspired by our results.
1 Introduction
Primary care providers are essential to the healthcare ecosystem because they are the first point of contact for many patients (Pearson and Raeke, 2000; Wu et al., 2022). Patients rely on primary care providers for routine checkups and referrals to specialists. Moreover, care continuity can instill trust and improve medication uptake rates and patient health (Wu et al., 2022).
Unfortunately, high provider turnover rates frequently lead to patients without an assigned provider (Reddy et al., 2015). Provider turnover disrupts patient care and leads to worse care (Reddy et al., 2015). In principle, healthcare administrators reassign unmatched patients to other providers; however, in practice, the process takes months due to provider scarcity and the logistical burden of rematching and coordinating patient matches (Hedden et al., 2021). While many patients find their new provider quickly, others have to wait years to find a new provider due to large numbers of patients, high turnover rates, and provider scarcity (Hedden et al., 2021; Shanafelt et al., 2012).
Algorithms that automatically match patients and providers can reduce logistical hassle but require balancing patient autonomy and system-wide utility. For example, while automatically assigning each patient to a provider would decrease wait times, it also reduces patient autonomy because patients cannot select their provider (Entwistle et al., 2010; Gaynor et al., 2016). In contrast, offering patients full autonomy could lead to suboptimal matches; for example, patients who rematch quickly might prevent better matches for late-responding patients. In addition, granting patients full independence can be overwhelming and delay decisions (Bate and Robert, 2005). We propose to match patients and providers through an assortment framework (Shi, 2016; Rios and Torrico, 2023; Davis et al., 2013). Assortments are menus of options given to customers who select an option from such a menu (Bertsimas and Mišic, 2015). Assortments are used in e-commerce settings where platforms give customers a suggested list of products and have them select from this list (Tian et al., 2024). In an assortment setting, each patient receives a menu of providers and selects a provider from such a menu. These curated offer sets allow administrators to balance patient autonomy and system-wide match utility.
We devise policies for offering assortments in a patient-provider matching scenario. To do so, we first develop a model of patient-provider matching then construct a set of policies for offering assortments. We characterize how policy performance depends on problem characteristics such as patient-to-provider ratio, and we empirically detail the regimes under which each of our policies performs best. We then evaluate our policies in a real-world scenario constructed on top of a Medicare dataset to assess the performance of our policies in a real-world example.  [^1] We include all code at https://github.com/naveenr414/patient-provider
Our key question is the following: How should healthcare administrators design provider menus for patients to optimize system-wide match rates and quality?
1.1 Contributions
We develop a model of patient-provider matching using assortment offerings and analyze its impact on match rate and match quality. We characterize well-performing policies when varying properties such as the patient/provider ratio. We then empirically analyze these policies in a synthetic and real-world dataset built using a Medicare dataset. We use the insights from these studies to propose recommendations for designing patient-provider matching systems in practice.
Developing a Model of Patient-Provider Matching
We develop a model of patient-provider matching where administrators offer assortments upfront, then patients sequentially respond in a random order and select providers one by one. We do so because it captures two critical elements: 1) patients have autonomy in selecting providers and 2) administrators have some control over the matching process by varying assortments offered to patients. We model patient decisions through a choice model that captures how the set of providers offered impacts the final provider chosen. The selection of choice model is key to the fidelity of our model, so we analyze various choice models, including the uniform choice model and the multinomial logit choice model. Finally, we capture patient heterogeneity through a match quality matrix, which captures differences in match quality between pairs of patients and providers. Match quality can encompass geographical proximity, provider specialties, and concordance along language, race, and gender (Greenwood et al., 2020, 2018; Manson, 1988). Patients tend to match with higher quality providers, though the exact match probabilities are dictated by their choice model.
Constructing Assortment Policies
We characterize the optimal policy in the single-provider scenario and use these insights to construct a set of policies for assortment optimization. We first construct a baseline policy, greedy, which offers every provider to every patient. We demonstrate that the greedy policy can do arbitrarily poorly, so we build upon this by developing three new policies that adaptively vary assortment sizes: pairwise, group-based, and gradient descent. The pairwise policy offers each patient at most one provider, and we demonstrate an instance-dependent approximation guarantee for this policy. We then develop the group-based policy by augmenting the pairwise policy so we offer assortments to groups of patients. We demonstrate that such group-based maintains the match rate of the pairwise policy while potentially improving the match quality. For the gradient descent policy, we construct an objective function that provably lower bounds the match quality, and we use gradient descent to optimize this quantity.
Characterizing Policy Performance
We analyze the performance of our policies in a synthetic setting and empirically characterize the conditions under which each policy performs best. We demonstrate that the best-performing policy depends on two factors: the match probability and the ratio of patients to providers. When patients outnumber providers, as is common in many real-world health systems, the gradient descent policy performs best because the objective function is closer to the lower bound. When patient and provider counts are balanced, we show that the greedy policy achieves a high match quality when patients have low match probabilities, while the group-based policy achieves a high match quality when patients have high match probabilities.
We evaluate policies in a real-world scenario constructed via Medicare data (Medicare & Medicaid, 2025) and we show here that the gradient descent policy performs best. We extend this scenario to real-world considerations such as fairness and cognitive overload.
Contribution Summary
Overall, we make three contributions to the patient-provider matching and assortment optimization literature: we i) develop a model of patient-provider matching under assortments, ii) characterize the performance of assortment policies, and detail how the best-performing policy depends on problem characteristics such as patient match probability and patient/provider ratio, and iii) make recommendations for the implementation of a real-world patient-provider matching system based on the results of our theoretical and empirical analysis.
Figure 1: A) In this work, we match patients with providers B) by offering a set of assortments to patients, where each assortment consists of a set of providers. C) Patients then respond sequentially and either select a provider from their assortments or abstain. D) This results in matches between patients and providers.
1.2 Related Work
Matching Patients and Providers Care continuity is critical for patient health, as it allows for better patient communication, lowers operating costs, and improves provider teamwork (Plomondon et al., 2007). Despite this, provider turnover rates are high, with the average healthcare system reporting a 7% turnover rate year-to-year (Plomondon et al., 2007). Moreover, such issues are exacerbated recently due to high rates of provider burnout (Shanafelt et al., 2012) and the Covid-19 pandemic (Shanafelt et al., 2022). Once a provider drops out, it is difficult for patients to find a new one; while 54% of patients find their new provider within 12 months, 6% fail to find a new provider even after 36 months (Hedden et al., 2021). High levels of provider turnover can worsen primary care because patients lack access to routine checkups or reminders to take medication (Reddy et al., 2015; Levine et al., 2019; Nguyen et al., 2020; Alsan et al., 2019).
While provider turnover impacts patient satisfaction, quantifying “match quality” is hard. Patients care mainly about geographic proximity (Dunlea and Lenert, 2015), with concordance along race (Greenwood et al., 2020), gender (Greenwood et al., 2018), and language (Manson, 1988) playing a secondary role. Communication style alignment can improve match quality (Street et al., 2008). These factors build trust and lead patients to more truthfully report symptoms (Persky et al., 2013).
Algorithms for Patient-Provider Matching Existing work in patient-provider matching uses either a one-shot matching framework or a genetic programming-based approach. Within the one-shot framework, linear programs can manage provider panel sizes (Harrington et al., 2021), the deferred acceptance framework can perform two-stage matching (Chen et al., 2020), and scheduling algorithms are used to model optimal rates of on-demand scheduling (Qu et al., 2007).
A second approach to patient-provider matching builds matches through genetic programming. Genetic programming algorithms learn matches over time through a fitness function. Within this framework, one line of work uses genetic programming to find fair matches between patients and providers (Chen et al., 2020), while another approach uses genetic programming to balance workloads between providers (Zhu et al., 2023). We build on these lines of work by incorporating patient autonomy into the process, which motivates the need for an assortment-based approach.
Assortment Optimization Our work analyzes patient-provider matching through assortment optimization, where retailers offer a set of menus to customers. In assortment optimization, retailers construct assortments and customers make decisions based on offered options (Bertsimas and Mišic, 2015). Retailers first construct models of customer decisions, such as the multinomial logit model (Davis et al., 2013) and nested logit models (Alfandari et al., 2021), both of which use logit probabilities to model customer decisions. Assortments are employed in domains including school choice (Shi, 2016), dating markets (Rios and Torrico, 2023), and e-commerce (Tian et al., 2024).
Assortment optimization algorithms come in offline and online variants. In offline assortment optimization, assortments are offered one-shot and customers make decisions simultaneously and independently (Bertsimas and Mišic, 2015). In online assortment optimization, customers arrive sequentially, and options are offered for each (Aouad and Saban, 2023). Our setting resides between offline and online, as we offer assortments offline, but patients respond online.
Matching under Random Ordering Patient-provider matching is challenging from an optimization perspective because patients make decisions in a random order. Randomized response orders are found in variants of online selection problems such as secretary problems (Freeman, 1983) and online learning (Orabona, 2019). One key difference is the lack of flexibility in our scenario; changing assortments online for the patient-provider matching setting would involve considerable logistical hassle, so we offer assortments in a one-shot manner.
2 Problem Formulation
2.1 Formal Model of Patient-Provider Matching
We focus on matching $N$ patients to $M$ providers. We present each patient $i$ with an assortment $\mathbf{X}_{i}\in\{0,1\}^{M}$ upfront. Here, $X_{i,j}$ denotes whether patient $i$ has provider $j$ in their assortment. Patients then select providers from their assortment sequentially by responding in an order $\sigma$, and can abstain from provider selection. Here, $\sigma_{1}\in[N]$ represents the first patient in the order. We assume patient response times are i.i.d. and exponential; prior work demonstrates this can model real-world wait times (Newell, 2013). Under this assumption, $\sigma$ is a random permutation of $[N]$.
Each patient selects a provider from their assortment according to a choice model $f_{i}:\{0,1\}^{M}\rightarrow\{0,1\}^{M}$. The choice model takes as input a 0-1 vector $\mathbf{z}$ and outputs a random 0-1 vector representing the provider selected by patient $i$. Formally, let $\mathbf{y}^{(t)}$ represent the set of unselected providers at time $t$. Then, patient $\sigma_{t}$ selects providers according to $f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})$, where $\odot$ is the element-wise product ($\mathbf{a}\odot\mathbf{b}=\mathbf{c}\rightarrow a_{k}b_{k}=c_{k}\forall k$). $\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}$ represents the set of providers that are on patient $\sigma_{t}$’s assortment and unselected by previous patients. If $f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})=\mathbf{0}$, then no provider is selected, while if $f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})_{j}=1$, then provider $j$ is selected. We note that a patient can match with at most one provider: $\lVert f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})\rVert_{1}\leq 1$
Patients are more likely to select providers with a higher match quality, $\theta_{i,j}\in[0,1]$, where $\theta_{i,j}$ is analogous to a reward for patient $i$ when selecting provider $j$. Match quality can encompass various factors, including demographic concordance (Greenwood et al., 2020, 2018; Manson, 1988), physical proximity, and patient needs. We assume administrators know $\theta_{i,j}$. Prior work demonstrates providers are able to successfully identify new high-match quality providers for patients (Lohr et al., 2013), which implies that we can predict match quality from patient data.
We present examples of choice models below:
  1. 1.
    Uniform - Patients select their most preferred provider with probability $p$ and otherwise abstain. Let $\mathbf{e}_{k}$ denote the $k$-th standard basis vector, then:
    $f_{i}(\mathbf{z})=\left\{\begin{array}[]{ll}\mathbf{e}_{\operatorname*{arg\,max}_{j}(\theta_{i}\odot\mathbf{z})_{j}}&\text{with prb. p}\\ \mathbf{0}&\text{otherwise}\end{array}\right.$ (1)
  2. 2.
    Threshold - Patients follow the uniform choice model but only select providers with match quality above some threshold $\alpha$. If match quality corresponds to geographic proximity, $\alpha$ corresponds to maximum travel distance. Formally:
    $f_{i}(\mathbf{z})=\left\{\begin{array}[]{ll}\mathbf{e}_{\operatorname*{arg\,max}_{j}(\theta_{i}\odot\mathbf{z})_{j}}&\text{with prb. p if }\max\theta_{i}\odot\mathbf{z}\geq\alpha\\ \mathbf{0}&\text{otherwise}\end{array}\right.$ (2)
  3. 3.
    Multinomial Logit (MNL) - Patients select providers according to $\theta$, and have an exit option $\gamma$ which represents selecting no provider. Under the MNL choice model (Train, 2009), this is:
    $f_{i}(\mathbf{z})=\left\{\begin{array}[]{ll}\mathbf{e}_{j}&\text{with prb. }\frac{z_{j}\exp(\theta_{i,j})}{\exp(\gamma)+\sum_{j^{\prime}}z_{j^{\prime}}\exp(\theta_{i,j^{\prime}})}\\ \\ \mathbf{0}&\text{with prb. }\frac{\exp(\gamma)}{\exp(\gamma)+\sum_{j^{\prime}}z_{j^{\prime}}\exp(\theta_{i,j^{\prime}})}\end{array}\right.$ (3)
We aim to find a policy $\pi(\theta)$ that constructs an assortment $\mathbf{X}$, given the match quality matrix $\theta$. We evaluate policies through two metrics: match quality and match rate.
  1. 1.
    Match Quality measures $\theta$ across selected patient-provider pairs. Formally, it aggregates $\theta_{\sigma_{t}}$ across provider selections $f_{\sigma_{t}}(\pi(\theta)\odot\mathbf{y}^{(t)})$ for all $N$ patients (= $N$ time periods):
    $\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}(\pi(\theta)\odot\mathbf{y}^{(t)})\cdot\theta_{\sigma_{t}}]$ (4)
  2. 2.
    Match Rate is the proportion of patients who are matched. Formally, it computes if patient $\sigma_{t}$ matches with a provider: $\lVert f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})\rVert_{1}\geq 0$:
    $\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}\lVert f_{\sigma_{t}}(\pi(\theta)\odot\mathbf{y}^{(t)})\rVert_{1}]$ (5)
Example 1.
We quantify the example in Figure 1. Suppose $N=M=3$, and let $\theta$ and $\mathbf{X}$ be
$\theta=\begin{bmatrix}0.5&0.6&0.9\\ 0.7&0.5&0.4\\ 0.6&0.2&0.8\end{bmatrix},\hskip 14.22636pt\mathbf{X}=\begin{bmatrix}0&0&1\\ 1&1&0\\ 1&0&1\end{bmatrix}$ (6)
Suppose $\sigma=[2,1,3]$. Then one trial could be the following: patient $\sigma_{1}=2$ abstains from selecting a provider, patient $1$ selects provider $3$ and patient $3$ selects provider $1$: $f_{2}(\mathbf{y}^{(1)}\odot\mathbf{X}_{2})=\mathbf{0}$, $f_{1}(\mathbf{y}^{(2)}\odot\mathbf{X}_{1})=\mathbf{e}_{3}$, and $f_{3}(\mathbf{y}^{(3)}\odot\mathbf{X}_{3})=\mathbf{e}_{1}$. Here, $\mathbf{y}^{(1)}=[1,1,1]$, $\mathbf{y}^{(2)}=[1,1,1]$, and $\mathbf{y}^{(3)}=[1,1,0]$. The resulting match quality is $\frac{1}{3}(\theta_{1,3}+\theta_{3,1})=0.5$ and match rate is $\frac{2}{3}$
2.2 Model Fidelity
We discuss our model fidelity for real-world patient-provider matching:
  1. 1.
    Patient autonomy is a critical part of our model, as assortments allow patients to select between providers while allowing administrators some control. Prior work demonstrated that people are willing to trade autonomy for efficiency, including in online markets (Diederich et al., 2025) and domestic life (Munro et al., 2014). We find similar ideas during our conversations with our healthcare partners, where patients might prefer a few high-quality choices to an overwhelmingly large number of options (Chernev et al., 2015).
  2. 2.
    Match quality and match rate underlie assortment computation. We focus on match quality because it reflects the benefit patients get from matching, while match rate reflects the fraction of patients with access to providers. For example, if patients primarily care about proximity, match quality should be based on the distance between patients and providers. We discuss different notions of match quality in Section 5.3.1.
  3. 3.
    Choice models reflect how patients select providers. Choice models are critical to our setting, so we analyze various choice models in Section 3 and experiment with different selections in Section 5.2. We focus on choice models that are clinically motivated or ubiquitous in prior work. For example, the uniform choice model reflects low-effort patient decisions (Salisbury, 1989), while the MNL choice model is commonly seen in prior work (Davis et al., 2013).
2.3 Analysis in the Single Provider Scenario
To gain insight into the optimal match quality policy structure, we analyze the $M=1$ scenario under the uniform choice model. We first explicitly detail the optimal policy: {theorem}[] Let $f_{i}$ be the uniform choice model with probability $p$. Suppose $M=1$, and let $u_{1},u_{2},\ldots,u_{N}$ be a set of coefficients such that $\theta_{u_{1}}\geq\theta_{u_{2}}\cdots\theta_{u_{N}}$. Let $s$ be defined as follows:
$s=\operatorname*{arg\,max}_{s}(1-(1-p)^{s})\frac{\sum_{i=1}^{N}\theta_{u_{i},1}}{s}$ (7)
Then the policy which maximizes match quality is $\mathbf{X}_{u_{1},1}=\mathbf{X}_{u_{2},1}\cdots\mathbf{X}_{u_{s},1}=1$, where $\mathbf{X}_{i,1}$ is 0 otherwise. When $M=1$, we can decompose the match quality into the probability any patient matches and the average match quality given matching. If we offer the single provider to $s$ patients, then the match probability is $(1-(1-p)^{s})$. For fixed $s$, we offer provider $j=1$ to the top-$s$ patients by match quality. Note that the optimal match rate is achieved with $s=N$, while the optimal match quality might arise from $s<N$ (as we’ll show in Section 3.5). We include all full proofs in Appendix B.
3 Constructing Matching Policies
We introduce policies $\pi(\theta)$ to construct assortments. We first introduce a baseline greedy policy and discuss why such a policy can perform poorly. We then propose three new policies: pairwise, which assigns each provider a single patient, group-based, which improves upon the weaknesses of the pairwise policy, and gradient descent, which optimizes for a lower bound on the match quality.
3.1 Greedy Policy
The greedy policy offers all providers to each patient and grants patients full autonomy. Formally, the greedy policy is $\pi^{R}(\theta)_{i,j}=1$, for every patient $i$ and provider $j$. The greedy policy maximizes the match rate under the uniform choice model because $\mathbf{X}_{\sigma_{t}}=\mathbf{1}$, and so $\lVert f_{\sigma_{t}}(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)})\rVert_{1}=\lVert f_{\sigma_{t}}(\mathbf{y}^{(t)})\rVert_{1}$.
For match quality, the greedy policy considers assortments myopically and ignores the structure of the match quality matrix, which can lead to poor performance. For example, if $M=1$ and $\theta_{1,1}>>\theta_{1,2}$, the greedy policy gives both patients equal chances of matching, whereas the optimal policy offers $j=1$ only to patient $i=1$. Formally, let $\pi^{*}$ be the optimal policy, $\pi^{*}(\theta)=\operatorname*{arg\,max}_{\mathbf{X}}\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$. We first prove that $\pi^{R}$ is an $\epsilon$-approximation:
{theorem}
[] Let $f_{i}$ be the uniform choice model with match probability $p$. For any $p$ and $\epsilon$, there exists a $\theta$ such that
$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{R}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]\leq\epsilon\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (8)
We prove this by generalizing the example where $\theta_{1,1}>>\theta_{2,1}$ to arbitrary $N$ and $M$. Here, the optimal policy is to offer each provider to only one patient. However, the greedy policy offers all providers to all patients, which leads to sub-optimal matches for each provider $j$. Moreover, we show that we can construct scenarios with $N=M=\frac{3}{\epsilon p}$ where the greedy policy is an $\epsilon$-approximation. For large $p$, even small $M$ and $N$ lead to situations where the greedy policy performs poorly.
3.2 Pairwise Policy
The pairwise policy inverts the greedy policy by offering each patient at most one provider. This strategy reduces patient choices while better incorporating match quality structure. The pairwise policy, $\pi^{P}(\theta)$, pairs patients and providers by solving the weighted bipartite matching problem:
$\max\limits_{\pi^{P}(\theta),\sum_{i=1}^{N}\pi^{P}(\theta)_{i,j}\leq 1,\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\leq 1}\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}$ (9)
Under a uniform choice model, the choice of $p$ impacts the pairwise policy performance: {theorem}[] Let $f_{i}$ be the uniform choice model with match probability $p$. Then
$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{P}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]\geq p\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (10)
To prove this, we first upper bound the optimal policy, $\pi^{*}(\theta)$, with the value of the underlying bipartite matching problem. We then show that the pairwise policy corresponds to the bipartite matching problem, where edges correspond to pairs of patients and providers, and each edge exists with probability $p$. Summing over edges gives that the pairwise policy is a $p$-approximation to the bipartite matching problem and is, therefore, a $p$-approximation to $\pi^{*}$.
Beyond the uniform choice model, we can show a similar guarantee in the MNL scenario. For the MNL choice model, the exit option $\gamma$ is analogous to $p$ in controlling policy performance. First, let $v_{i}=j$ if $\pi^{P}(\theta)_{i,j}=1$ and let $v_{i}=-1$ if $\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}=0$. Then the MNL choice model is equivalent to the uniform choice model with $p_{i}=\frac{\mathbbm{1}[v_{i}\geq 0]\exp(\theta_{i,v_{i}})}{\exp(\gamma)+\mathbbm{1}[v_{i}\geq 0]\exp(\theta_{i,v_{i}})}$. We can then upper bound optimal match quality as $\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}[v_{i}\geq 0]\theta_{i,v_{i}}$, while the pairwise policy achieves $\sum_{i=1}^{N}\frac{p_{i}}{N}\mathbbm{1}[v_{i}\geq 0]\theta_{i,v_{i}}$. Let $a_{i}=\mathbbm{1}[v_{i}\geq 0]\theta_{i,v_{i}}$, then the pairwise policy achieves an approximation ratio of:
$\frac{\sum_{i=1}^{N}a_{i}\frac{\exp(a_{i})}{\exp(\gamma)+\exp(a_{i})}}{\sum_{i=1}^{N}a_{i}}\geq\frac{\exp(\frac{1}{N}\sum_{i=1}^{N}a_{i})(\frac{1}{N}\sum_{i=1}^{N}a_{i})}{(\exp(\gamma)+\exp(\frac{1}{N}\sum_{i=1}^{N}a_{i}))(\frac{1}{N}\sum_{i=1}^{N}a_{i})}\geq\frac{\exp(\frac{1}{N}\sum_{i=1}^{N}a_{i})}{\exp(\gamma)+\exp(\frac{1}{N}\sum_{i=1}^{N}a_{i})}$ (11)
Small $\gamma$ improves pairwise performance, corresponding to patients more inclined to match.
3.3 Group-Based Policy
The pairwise policy performs poorly for small $p$, so we develop the group-based policy. When $p$ is small, administrators should offer patients larger assortments because a patient’s top preference is more likely to be available. For example, consider a scenario with $N$ patients and $M$ providers. We demonstrate a performance gap between greedy and pairwise policy dependent on $p$:
Proposition 3.1.
Let $f_{i}$ be the uniform choice model with match probability $p$. If $\theta_{i,j}\sim U(0,1)$ and $M\leq N$ then
$\frac{\mathbb{E}_{\sigma,\theta}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{R}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]}{\mathbb{E}_{\sigma,\theta}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{P}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]}\geq\frac{1-(1-p)^{N/M}}{2p}$ (12)
We demonstrate this by showing that the pairwise policy achieves a match quality of $Mp$. We then compute the match probability for the greedy policy as $1-(1-p)^{N/M}$ for each of the $M$ providers, and noting that the match quality for each matched provider is at least $\frac{1}{2}$ in expectation. The pairwise policy performs poorly for small $p$, so we develop a new policy that modulates between the pairwise policy (for large $p$) and the greedy policy (for small $p$). To do this, we add a provider $j^{\prime}$ to the assortment of patient $i$ if $\theta_{i,j^{\prime}}\geq\theta_{i,v_{i}}$; that is, patients should be assigned preferred providers. However, such a policy could decrease match rates because some patients have empty assortments.
Example 2.
Let $\theta$ be the following:
$\theta=\begin{bmatrix}0.6&0.7\\ 0.3&0.6\\ \end{bmatrix}$ (13)
Here, $\pi^{P}(\theta)_{1,1}=\pi^{P}(\theta)_{2,2}=1$. We add provider $j=2$ to patient $i=1$, so $X_{i,2}=1$. Let $\sigma=[1,2]$. Then, with probability $p$, provider $j=2$ is selected by patient $i=1$ and $\mathbf{y}^{(2)}=[1,0]$ so patient $i=2$ has no options. This reduces the match rate from $p$ to $p-\frac{p^{2}}{2}$.
We overcome this issue by developing the group-based policy, $\pi^{G}(\theta)$. The group-based policy constructs groups so that each group receives the same assortment. In Example 2, this would mean that $\mathbf{X}_{1}=\mathbf{X}_{2}=[1,1]$. We demonstrate that such a grouping strategy is the only way to maintain a match rate while building upon the pairwise policy: {theorem}[] Let $\pi$ be a policy that augments the pairwise policy; $\pi(\theta)_{i,j}\geq\pi^{P}(\theta)_{i,j}$ for any $\theta$ for all $i,j$. Let $G=(V,E)$ be a directed graph with $N$ nodes such that nodes $i$ and $i^{\prime}$ are connected if $\pi(\theta)_{i,v(\theta)_{i^{\prime}}}=1$ Here, $v(\theta)_{i}=j$ if $\pi^{P}(\theta)_{i,j}=1$. If $N=M$, then
$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}\lVert f_{\sigma_{t}}\left(\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\rVert_{1}]=\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}\lVert f_{\sigma_{t}}\left(\pi^{P}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\rVert_{1}]=p$ (14)
if and only if each component in $G$ is a complete digraph. We prove this by showing that when $N=M$, all patients can match with a provider with probability $p$, leading to an overall match rate of $p$. To construct groups, we first compute edge weights for each pair of patients in $G$, denoted as $\alpha_{i,i^{\prime}}$. Here, $\alpha_{i,i^{\prime}}$ captures the benefit when placing $(i,v_{i})$ and $(i^{\prime},v_{i^{\prime}})$ in the same assortment. Next, we construct groups to maximize the sum of pairwise edges within each group. We do so through a linear program, with $z_{i,j}$ representing whether we include edge $(i,j)$ and $q_{i}$ representing whether we include node $i$ in the group. We place details on computing $\alpha_{i,i^{\prime}}$ and constructing the assortments in Algorithm 1.
Algorithm 1 Group-based policy ($\pi^{G}$)
1:  Input: Pairwise policy, $\pi^{P}$ and match quality $\theta$
2:  Output: Assortment, $\mathbf{X}$
3:  Initialize $\mathbf{X}=\mathbf{X}^{P}=\pi^{P}(\theta)$, $v_{i}=j$ if $X_{i,j}^{P}=1$, and $\mathbf{c}=\mathbf{1}$
4:  for all $(i,i^{\prime})\subseteq[N]$ do
5:     Let $\mathbf{X}^{\prime}=\mathbf{X}^{P}$
6:     Let ${X}^{\prime}_{i,v_{i^{\prime}}}=1$ and $X^{\prime}_{i^{\prime},v_{i}}=1$
7:     Let $\alpha_{i,i^{\prime}}=\mathbb{E}_{\sigma\sim[\{i,i^{\prime}\},\{i^{\prime},i\}]}[\sum_{t=1}^{2}(f_{\sigma_{t}}\left(\mathbf{X}^{\prime}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}-f_{\sigma_{t}}\left(\mathbf{X}^{P}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$
8:  end for
9:  Let $s=\infty$ and $r=\{1,\ldots,N\}$
10:  while s ¿ 0 do
11:     Let $s=\max\limits_{\mathbf{z}}\sum_{i\in r}\sum_{i^{\prime}\in r}z_{i,j}\alpha_{i,i^{\prime}}$, subject to $z_{i,j}=q_{i}q_{j}$, and let $\mathbf{z}$ be the corresponding solution
12:     for $i,i^{\prime}\in r$ with $z_{i,i^{\prime}}=1$ do
13:        Let $X_{i,v_{i^{\prime}}}=X_{i^{\prime},v_{i}}=1$
14:        Remove $i,i^{\prime}$ from $r$
15:     end for
16:  end while
3.4 Gradient Descent Policy
The group-based policy makes local improvements to the pairwise policy but can fail because it only considers pairwise interactions. To fix this, we develop the gradient descent policy by considering the global structure in $\theta$. Our gradient descent policy uses gradient descent to optimize an objective function $g(f(\mathbf{X}))$, which represents a lower bound on the match quality of an assortment $\mathbf{X}$ under the uniform choice model. Here, $f(\mathbf{X})_{i,j}$ lower bounds the probability that provider $j$ is available, while $g(f(\mathbf{X}))_{i,j}\in[0,1]^{N\times M}$ lower bounds the probability that provider $j$ is the top choice for patient $i$: $g(f(\mathbf{X}))\leq\mathrm{Pr}[j=\operatorname*{arg\,max}_{j^{\prime}}\theta_{i,j^{\prime}}\mathbf{y}^{(t)}_{j^{\prime}}]$. We first detail how to compute $f$ and $g$, then prove that this lower bounds the match quality (see Algorithm 2):
  1. 1.
    $f(\mathbf{X})$ lower bounds the availability probability for provider $j$. We first note that $f(\mathbf{X})_{i,j}=0$ if $X_{i,j}=0$. Next, availability depends on the response order $\sigma$. For example, if $\sigma_{1}=i$, then provider $j$ is available if $X_{i,j}=1$. Next, if $\sigma_{t}=i$, then under a uniform choice model, there is at least a $(1-p)^{t-1}$ availability probability. This holds because, in the worst case, the $t-1$ preceding patients also prefer provider $j$, so the probability none of the $t-1$ patients select $j$ is $(1-p)^{t-1}$. Next, if $n$ out of $N$ patients have provider $j$ on their assortment, then the probability that a given patient $i^{\prime}\neq i$ matches to $j$ is $p\frac{n-1}{N-1}$. That is, if patient $i$ is in the t-th position, then the availability probability is at least $(1-p\frac{n-1}{N-1})^{t-1}$. Averaging over all $N$ potential values for $t$ gives:
    $f(\mathbf{X})_{i,j}=X_{i,j}h(\lVert\mathbf{X}_{*,j}\rVert_{1}),h(n)=\frac{1}{N}\sum_{t=1}^{N}(1-p\frac{n-1}{N-1})^{t-1}$ (15)
  2. 2.
    $g(f(\mathbf{X}))$ computes the probability that provider $j$ is the top available provider. Let $u_{i,1},u_{i,2},\ldots,u_{i,N}$ be coefficients so $\theta_{i,u_{1}}\geq\theta_{i,u_{2}}\cdots\theta_{i,u_{M}}$. Then provider $u_{i,1}$ is the top option whenever it is available; $g(f(\mathbf{X}))_{i,u_{i,1}})=f(\mathbf{X})_{i,u_{i,1}}$. Similarly, $u_{i,2}$ is only selected if $u_{i,1}$ is not selected; we can estimate $g(f(\mathbf{X}))_{i,u_{i,2}})=f(\mathbf{X})_{i,u_{i,2}}(1-f(\mathbf{X})_{i,u_{i,1}})$. We can generalize this as:
    $g(f(\mathbf{X}))_{i,u_{i,k}}=f(\mathbf{X})_{i,u_{i,k}}\prod_{k^{\prime}=1}^{k-1}(1-f(\mathbf{X})_{i,u_{i,k^{\prime}}})$ (16)
    The expected match quality is then $p(g(f(\mathbf{X}))\cdot\theta)$.
Algorithm 2 Gradient Descent Objective
  Input: Match quality $\theta$ and Assortment $\mathbf{X}$
  Output: Estimated total match quality
  Let $h(n)=\frac{1}{N}\sum_{t=1}^{N}(1-p\frac{n-1}{N-1})^{t-1}$
  Let $f(\mathbf{X})_{i,j}=h(\lVert\mathbf{X}_{*,j}\rVert_{1})\mathbf{X}_{i,j}$
  Let $u_{i,1},u_{i,2},\ldots,u_{i,M}$ be a permutation of $[M]$, so $\theta_{i,u_{i,1}}\geq\theta_{i,u_{i,1}}\cdots\theta_{i,u_{i,M}}$
  Let $g(f(\mathbf{X}))_{i,u_{k}}=f(\mathbf{X})_{i,u_{k}}\prod_{k^{\prime}=1}^{k-1}(1-f(\mathbf{X})_{i,u_{k^{\prime}}})$
  return  $p*g(f(\mathbf{X}))\cdot\theta$
{theorem}
The following holds for any $\mathbf{X}$:
$p*(g(f(\mathbf{X}))\cdot\theta)\leq\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (17)
Moreover, we demonstrate tightness when using the pairwise policy: {lemma}[] When $\mathbf{X}=\pi^{P}(\theta)$, then
$p*(g(f(\mathbf{X}))\cdot\theta)=\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (18)
We demonstrate this by showing that for any $i$, $\sum_{j=1}^{m}pg(f(\mathbf{X}))_{i,j}$ is an underestimation for the probability that any of the top-$m$ providers are available. We then show that because $\theta_{i,u_{i,j}}$ is sorted from high to low, we can similarly prove that $\sum_{j=1}^{m}p(g(f(\mathbf{X}))_{i,j}\theta_{i,j}$ is also an underestimation for each patient $i$. In the scenario where $\mathbf{X}=\pi^{P}(\theta)$, we can explicitly compute $g(f(\mathbf{X}))$ because each patient matches with at most one provider.
Key to our policy is the tightness of the lower bound. When each patient receives more providers the lower bound is looser so the gradient descent policy performs worse. Conversely, when $N>>M$, each patient can only receive a few providers, so our policy performs better. Essentially, when there is low “inter-provider interference”, or patients receiving large assortments (relative to the number of providers), our policy performs better. The gradient descent policy can also naturally extend to optimize match rate if $\theta_{i,j}$ is constant; in that scenario, we optimize for $\sum_{i=1}^{N}\sum_{j=1}^{M}g(f(\mathbf{X}))_{i,j}$.
3.5 Worked Example with Policies
We motivate the need for different policies beyond greedy through a simple example with $N=3$ patients and $M=1$ providers.
Example 3.
Consider a scenario with $p=0.75$ and $\theta=[0.7,0.7,0.1]$. According to Theorem 2.3, the optimal assortment in such a scenario is $\mathbf{X}=[1,1,0]$, which achieves a match quality of $0.22$. We next compare our four policies:
  1. 1.
    Greedy offers $\mathbf{X}=[1,1,1]$, which leads to an average match quality of $\frac{(1-(1-p)^{N})\frac{1}{N}\sum_{i=1}^{N}\theta_{i,1}}{N}=0.16$. Greedy achieves a low match quality because it offers a provider to patient $i=3$, which drags down the average match quality.
  2. 2.
    Pairwise offers $\mathbf{X}=[1,0,0]$, which achieves a match quality of $0.18$. While the pairwise policy improves upon the greedy policy, the symmetry between patients one and two implies that it should offer a provider to patient $i=2$, as it can improve the match probability.
  3. 3.
    Group-Based first computes $\alpha_{1,2}=0.13$ and $\alpha_{1,3}=-0.15$ and constructs a group with patients 1 and 2. It then offers $\mathbf{X}=[1,1,0]$, the optimal assortment for match quality.
  4. 4.
    Gradient Descent optimizes for $g(f(\mathbf{X}))\cdot\theta=f(\mathbf{X})\cdot\theta=\sum_{i=1}^{N}X_{i,1}\theta_{i,1}h(\lVert\mathbf{X}_{*,j}\rVert_{1})$, which is optimizes at $\mathbf{X}=[1,1,0]$ and is the optimal assortment for match quality.
The greedy policy fails because it offers the provider to patients with a low match quality, while the pairwise policy fails because it offers providers to too few patients. The group-based and gradient descent policies rectify this situation by adapting assortment size based on the match quality and correctly offering the provider to patients 1 and 2.
4 Extending Policies to Real-World Considerations
We extend our model to incorporate three real-world matching phenomena: 1) batch-offering assortments, 2) patient cognitive load, and 3) fairness considerations.
4.1 Batch Offering Patients
Thus far, we have assumed that all menus are constructed up front, in a single “batch.” Healthcare administrators can sacrifice logistical ease to improve match quality by making offers to groups of patients in batches rather than all at once. By offering assortments in batches, administrators increase control over the response order but increase the total response time for all patients to respond. By assumption, the response time for each patient is exponentially distributed (see Section 2.1), and the expected maximum of $N$ exponential variables is proportional to $\log(N)$ (Eisenberg, 2008). By offering $L$ batches, our total response time increases from $\log(N)$ to $L\log(N/L)$.
We aim to construct batches, $b_{1},b_{2},\ldots,b_{L}\subseteq[N]$, so that the resulting response order maximizes match quality. Formally, let $S(b_{1},b_{2},\ldots,b_{L})$ be a distribution over response orders so that any patient $i\in b_{k}$ is before any patient $i^{\prime}\in b_{k^{\prime}}$ with $k<k^{\prime}$. Next, let $\sigma^{*}$ be the optimal ordering for some fixed policy $\pi$: $\sigma^{*}=\operatorname*{arg\,max}_{\sigma}\frac{1}{N}\sum_{t=1}^{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi(\theta,\sigma)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})$. We aim to set $b_{1},b_{2},\ldots,b_{L}$ so that $\sigma\sim S(b_{1},b_{2},\ldots,b_{L})$ and $\sigma^{*}$ are similar.
To do so, we first characterize $\sigma^{*}$ for a policy $\pi^{A}(\theta,\sigma)$, which augments the pairwise policy using knowledge of $\sigma$, $\pi^{A}(\theta,\sigma)_{i,j}\geq\pi^{P}(\theta)_{i,j}\forall i,j$. Recall that $v_{i}=j$ if $\pi^{P}(\theta)_{i,j}=1$, and let $v^{-1}_{j}=i$ if $v_{i}=j$. Then, $\pi^{A}(\theta)_{i,j}=1$ if $\theta_{i,j}\geq\theta_{i,v_{i}}$ and $\sigma^{-1}_{i}\geq\sigma^{-1}_{v^{-1}_{j}}$, where $\sigma^{-1}_{i}=t$ if $\sigma_{t}=i$. Essentially, $\pi^{A}(\theta)$ augments the pairwise policy $\pi^{P}(\theta)$ to add pairs $i,v_{i^{\prime}}$ for patients $i^{\prime}$ preceding $i$ in $\sigma$. We focus on such a policy because it represents a natural extension of the pairwise policy to incorporate the order $\sigma$. We characterize $\sigma^{*}$ for the policy $\pi^{A}(\theta)$ then use this to construct the batches $b_{k}$:
Lemma 4.1.
Let $G=(V,E)$ be a graph with $N$ nodes such that node $i$ is connected to $i^{\prime}$ if $\theta_{i,v_{i}}\leq\theta_{i,v_{i^{\prime}}}$ or $v_{i}=-1$ and $v_{i^{\prime}}\neq-1$. Let $\sigma^{*}$ be the optimal ordering for the policy $\pi^{A}$. Then, traversing the nodes defined by $\sigma^{*}$ is a reverse topological ordering on $G$.
We demonstrate this by considering a pair of patients $i$ and $i^{\prime}$, so that $i$ precedes $i^{\prime}$ in any reverse topological order, but $i$ precedes $i^{\prime}$ in $\sigma^{*}$. We then use case work to show that we can either a) show that there is a cycle in the graph $G$, leading to a contradiction, b) recurse on a smaller pair of patients, $w,i$ or $i^{\prime},w$, or c) show that swapping $i$ and $i^{\prime}$ in $\sigma$ can only increase the match quality.
We next construct $b_{1},b_{2},\ldots,b_{L}$, so $\sigma\sim S(b_{1},b_{2},\ldots,b_{L})$ is close to the optimal ordering $\sigma^{*}$. The idea is to partition $\sigma^{*}$ into $L$ batches; then we know that all patients in $b_{1}$ precede $b_{2}$. To do this, we first demonstrate that if no pairs of patients within a batch share a common descendant, then such a batching preserves the match quality of $\sigma^{*}$: {theorem}[] Let $\sigma^{*}$ be an optimal ordering. Let $G=(V,E)$ be a graph with $N$ nodes such that node $i$ is connected to $i^{\prime}$ if $\theta_{i,v_{i}}\leq\theta_{i,v_{i^{\prime}}}$ or $v_{i}=-1$ and $v_{i^{\prime}}\neq-1$. Suppose that there exists a partition of $\sigma^{*}$ into $K$ batches, $b_{1},b_{2},\ldots,b_{L}$, such that for any partition $b_{k}$, no $i,i^{\prime}\in b_{k}$ have a common descendant in $G=(V,E)$. Then $\pi^{A}$ achieves the same match quality under partition $g$ as under the optimal ordering:
$\mathbb{E}_{\sigma\sim S(b_{1},b_{2},\ldots,b_{L})}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{A}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]=\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma^{*}_{t}}\left(\pi^{A}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})$ (19)
This holds because the selection made by patient $i$ depends only on the descendants of patient $i$ on the graph G. If $b_{1},b_{2},\ldots,b_{K}$ ensures that any pair of descendants are in different batches, then the order of descendants is maintained, so patients select the same providers in $\sigma$ and $\sigma^{*}$.
In practice, we compute the batches $b_{k}$ through the following:
  1. 1.
    Computing the graph $G=(V,E)$ with $i$ connected to $i^{\prime}$ if $\theta_{i,v_{i^{\prime}}}>\theta_{i,v_{i}}$.
  2. 2.
    Computing a reverse topological ordering $\sigma^{*}$.
  3. 3.
    Selecting batches with the minimum number of edges within a batch (via dynamic programming).
We note that this is an approximation of optimal batches for two reasons: 1) while $\sigma^{*}$ is a reverse topological ordering, there could be multiple reverse topological orderings, and 2) minimizing the number of edges within a batch does not guarantee that the order of descendants is maintained.
4.2 Maximum Assortment Sizes
To incorporate patient cognitive load (Chernev et al., 2015), we incorporate restrictions on the maximum assortment size into our model. Large assortments might be undesirable or practically infeasible because patients only consider a subset of the options they are presented due to logistical or time constraints. Formally, we suppose that patients only consider a random $d$-subset $\mathbf{z}^{(d)}$ of the options they are presented at random, $z^{(d)}_{j}\leq X_{i,j}y^{(t)}_{j}$ with $\lVert\mathbf{z}^{(d)}\rVert_{1}\leq d$. Patients make decisions according to $f_{i}(\mathbf{z}^{(d)})$ rather than $f_{i}(\pi(\theta)\odot\mathbf{y}^{(t)})$. When $d$ is small, policies should offer small tailored assortments, as large assortments lead to patients ignoring many options.
4.3 Fair Matches
Policies should aim to achieve fair matches to encourage participation and maintain equity. We construct examples of fairness metrics through the prior literature on fair matches and list these below (Xinying Chen and Hooker, 2023; García-Soriano and Bonchi, 2020).
  • Minimum Match Quality - One natural objective is to maximize the minimum match quality across all matched patients. Such an objective is natural when administrators aim to maintain a certain level of care across patients.
  • Variance in Match Quality - Another approach minimizes variance in match quality to ensure that the discrepancy in match quality between patients is small.
  • Match Quality Range - One final approach to match quality fairness is to minimize the absolute difference between the maximum and minimum match quality for matched patients.
Each notion of fairness has different implications for the optimal policy. For example, optimizing for the minimum match quality might lead policies to only offer assortments to high match quality patient-provider pairs, whereas minimizing the variance in match quality might yield many low quality matches. We explore tradeoffs in match quality and fairness in Section 5.4.
5 Empirical Analysis
We analyze policies in synthetic and real-world settings to characterize how assortment policy impacts match quality and rate. First, we demonstrate that gradient descent is the best policy when patients outnumber providers. We then show that group-based and greedy policies perform well for large and small values of match probability $p$, respectively. Our results demonstrate that assortment policies should be selected based on problem specifics.
5.1 Experimental Details
We compare the four policies from Section 3 (greedy, pairwise, group-based, and gradient descent) and a random baseline with $\pi(\theta)_{i,j}\sim\text{Ber}(\frac{1}{2})$. For a fixed problem setup, we randomly sample $T$ permutations of patients and compute the match quality and match rate for the corresponding patient order ($\sigma$). We average runs across 15 seeds and $T=100$ trials and plot the standard error across these seeds. Because the scale of match quality and rate might vary between experiments, we report the normalized match quality and match rate (norm. MQ and norm. MR for short), which divides each quantity by that of the random policy.
5.2 Synthetic Examples
We construct synthetic examples to gain insight into our policies across different scenarios. For each example, we specify values for $M$, $N$, $f_{i}$, and $\theta$.
Example 4.
Figure 2: We find the optimal policy and compute the approximation ratio for match quality compared to optimal. While all non-random policies achieve approximation ratios $\geq 0.9$ when $M\geq N$ (left), the pairwise policy performs poorly and achieves a low approximation ratio when $N\geq M$ (right).
Few patients and providers - We first analyze policies with a small number of patients $N$ and number of providers $M$. Small $N$ and $M$ allow us to compare to the optimal policy $\pi^{*}$ because we can compute $\pi^{*}$ by enumerating all $2^{NM}$ assortments and selecting the one that maximizes match quality. We run two experiments: the first fixes $N=2$ while varying $M$ from $2$ to $5$, and the second fixes $M=2$ while varying $N$ from $2$ to $5$. For both, we uniformly distribute $\theta$, $\theta_{i,j}\sim U(0,1)$, and let the choice model be uniform choice with $p=1/2$. We compute the match quality approximation ratio to understand how close each policy is to the optimal.
In Figure 2, we demonstrate that when $M\geq N$ (left), all non-random policies perform well and achieve within 3.5% of optimal, while for $N\geq M$ (right), the pairwise policy performs poorly. When $M\geq N$, providers are abundant so each patient can match with their most preferred provider. However, when $N\geq M$, the pairwise policy performs up to 28% worse than optimal. Providers are scarce in this scenario, so policies should offer the same provider to multiple patients. The pairwise policy only offers each provider one patient, which is suboptimal and leads to poor match quality.
Example 5.
Figure 3: When there are more patients than providers, the gradient descent policy performs best, both for uniformly and normally distributed $\theta$.
Varied patient/provider ratio - In real-world situations, patients outnumbers providers (HHS, 2013), so we analyze how the patient/provider ratio impacts policy performance. We fix the number of providers $M=25$, while varying the number of patients $N$ from $25$ up to $200$. We compare policies under two match quality distributions: the first uniformly distributes $\theta\sim U(0,1)$, while the second normally distributes $\theta_{i,j}\sim\mathcal{N}(\mu_{j},0.1)$. The former corresponds to heterogeneous preferences where patient preferences are non-correlated, while the latter corresponds to homogenous preferences, where patient preferences are influenced by some signal of provider quality $\mu_{j}$. For both experiments, we fix $p=0.5$ for the uniform choice model.
In Figure 3, we demonstrate that the gradient descent policy performs best when patients outnumber providers. When $N/M=8$, the gradient descent policy performs 5% better than alternatives for uniform $\theta$ and 9% for normal $\theta$. The gradient descent policy performs best for large $N/M$ because $g(f(\mathbf{X}))$ better approximates match quality when $N>M$; see Section 3.4 for details. Both greedy and group-based policies perform similarly, while the pairwise policy performs worst, mirroring trends from Example 4. In Appendix A, we similarly compare policies according to the match rate and find that the gradient descent and greedy policies both maximize the match rate.
Example 6.
Figure 4: The greedy policy performs poorly when $p=0.9$ compared to smaller $p$, for both uniform and normal $\theta$, while the pairwise policy performs worse for smaller $p$, especially when $\theta$ is normally distributed. The group-based policy gets around these issues and performs well when $p$ is large and small.
Impact of $p$ and $\theta$ We assess the impact of match quality $\theta$ and match probability $p$ on policy performance with balanced patient and provider counts. We fix $N=M=25$, vary $p\in\{0.1,0.25,0.5,0.75,0.9\}$, and distribute $\theta$ either uniformly or normally.
In Figure 4, we demonstrate that the group-based policy achieves high match quality. As noted in Proposition 3.1, the group-based policy can improve upon the pairwise policy when $p$ is small. The pairwise policy performs especially poorly when $\theta$ is normally distributed because patient preferences are homogeneous, as all patients prefer some provider $j^{*}=\operatorname*{arg\,max}_{j}\mu_{j}$. The pairwise policy only places $j^{*}$ in one patient’s assortment, whereas $j^{*}$ should optimally be in multiple assortments due to its preferability. We also note the greedy and group-based policies perform well across $p$ and $\theta$; both policies are always within 6% of the best-performing policy. In Appendix A, we demonstrate that policies achieve similar match rates across $\theta$ and $p$.
(a) Uniform
(b) Normal
Figure 5: We characterize which policy maximizes match quality when varying patient/provider ratio and match probability $p$ when $\theta$ is uniformly (left) and normally (right) distributed. The group-based policy performs best for large $p$, the greedy policy performs best for small $p$, and the gradient descent policy performs best when the patient/provider ratio is large.
To understand which policy performs best across $p$, $\theta$, and $N/M$, we plot the policy which maximizes match quality when varying $p\in\{0.1,0.3,0.5,0.7,0.9\}$ and $N/M\in\{1,2,3,4,5,6,7,8\}$ with $M=10$, and distributing $\theta$ uniformly (Figure 5a) and normally (Figure 5b). The gradient descent policy performs best for large patient/provider ratios, which matches the results from Example 5. The group-based policy performs well for small $p$ because the problem more closely resembles bipartite matching (see Theorem 9), while the greedy policy performs well for small $p$ because it becomes advantageous to offer larger assortments (see Proposition 3.1). We finally note the random policy performs best when $p=0.1$ and $N/M=1$ because policies have high variance in that situation, making it difficult to ascertain which policy performs best.
Example 7.
Varied Choice Model - We evaluate the impact of choice models on policy performance by investigating the MNL and threshold choice models. We fix $N=M=25$, let $\theta$ be distributed uniformly, and vary $\gamma\in\{0,0.1,0.25,0.5\}$ for the MNL choice model and vary $\alpha\in\{0,0.1,0.25,0.5,0.75\}$ with $p=0.5$ for the threshold choice model. .
Figure 6: We compare policy performances when varying the exit option $\gamma$ for the MNL choice model, and the threshold $\alpha$ for the threshold choice model. For the MNL choice model, when $\gamma$ is small, the greedy policy performs poorly, while for large $\gamma$, all non-random policies perform similarly. For the threshold choice model, all policies perform similarly across choices of $\alpha$.
In Figure 6, we show that $\gamma$ determines the performance of the greedy policy for the MNL choice model, while policies perform similarly across $\alpha$ for the threshold choice model. As noted in Section 3.2, the choice of $\gamma$ is inversely related to the setting of $p$, with small $\gamma$ analogous to large $p$. Similar to the trends in Figure 4, we find that the greedy policy performs 22% worse than the best policy when $\gamma=0.1$. For the threshold choice model, $\alpha$ has little impact on the best policy, as greedy, pairwise, and gradient descent perform similarly across $\alpha$. In Appendix A, we explore variants of the uniform choice model where the match probability $p$ is misspecified.
5.3 Real-World Dataset
We work with a partner healthcare organization to assess the performance of our policies in a simulation that mimics situations faced in practice.
5.3.1 Constructing Data from Healthcare Providers
We work with a large healthcare provider in Connecticut to develop a simulation that reflects their system dynamics. We select values of $N$, $M$, $\theta$, and $f_{i}$:
  1. 1.
    $N$ and $M$ We estimate $N=1225$ and $M=700$ from the average panel size and provider count of the organization.
  2. 2.
    Choice Model $f_{i}$ We select the choice model based on previous work which shows that patients are low-effort decision makers who primarily make decisions based on geographic proximity (Salisbury, 1989; Buzza et al., 2011). We incorporate these two factors and let $f_{i}$ be a threshold model, with $\alpha$ representing the maximum distance patients are willing to travel.
  3. 3.
    Match Quality $\theta$ We construct $\theta$ to reflect geographic proximity and the presence of comorbidities. Geographic proximity is the top factor impacting patient match quality (Yen and Mounts, 2013; Buzza et al., 2011), while patients with comorbidities are best served by providers with the corresponding specialized training. Formally, for a patient $i$ and provider $j$, let $d_{i,j}$ be the distance between the patient and provider, and let $\beta_{i,j}$ denote whether patient $i$’s comorbidity and $j$’s specialty match. For example, if patient $i$ has a heart-related comorbidity and provider $j$ has a cardiology speciality, then $\beta_{i,j}=1$. We manually match patient comorbidities to provider specialties. We therefore let match quality $\theta_{i,j}=\alpha+(1-\alpha)(\Delta\beta_{i,j}+(1-\Delta)(\frac{\bar{d}}{d_{i,j}}-1)$. Here, $\Delta$ weights between comorbidities and distance and $\alpha$ sets the threshold match quality for a patient at distance $d_{i,j}=\bar{d}$.
We compute $d_{i,j}$ by sampling patient locations from Connecticut zip codes and obtaining provider locations from a Medicare dataset (Medicare & Medicaid, 2025). We obtain $\beta_{i,j}$ from prior work on comorbidity rates  (Piccirillo et al., 2008) and using information on provider specialties from a Medicare dataset (Medicare & Medicaid, 2025). We let $p=0.75$ because most patients are low-effort, and $\Delta=\alpha=\frac{1}{2}$ to balance proximity and comorbidities. Finally, we set $\bar{d}=20.2$ since prior work demonstrates the average distance threshold for patients is 20.2 miles (Yen and Mounts, 2013).
5.3.2 Real-World Simulation Results
(a) Comparison
(b) Comorbidities
(c) Match Rate vs. Population
(d) Match Rate by Zipcode
Figure 7: (a) We construct a real-world scenario with $N=1225$ and $M=700$. We find that the gradient descent policy performs best because of the large patient/provider ratio. (b) The gradient descent policy achieves a higher match quality for patients with comorbidities because these patients match with specialist providers. (c) Geographically, the gradient descent policy tends to match patients from more populated zip codes, (d) with the highest match rates occurring in cities such as Hartford and New Haven.
In Figure 7a, we show that the gradient descent policy achieves the highest match quality and outperforms alternatives by 13%. This mirrors Figure 5, where gradient descent is the best policy for large patient/provider ratios. Next, we analyze which patients achieve high-quality matches under the gradient descent policy. We find that patients with comorbidities achieve higher match quality ($p<10^{-17}$) because they tend to match with specialist providers (Figure 7b). Additionally, patients from more populated zip codes achieve higher match rates (Figure 7c), potentially due to the lack of providers in rural areas (Hart et al., 2002). Pictorially, this corresponds to high match rates in urban areas such as Hartford and New Haven (Figure 7d).
5.4 Real-World Considerations
We extend our scenario to explore the real-world considerations introduced in Section 4: patient batching, maximum assortment sizes, and fairness considerations.
5.4.1 Batched Patients
(a) Batch Size
(b) Max Assortment Size
(c) Fairness
Figure 8: (a) Offering patients in two batches improves match quality by 7% while increasing response time by 90%. Beyond two batches, the increases in match quality are diminishing. (b) We simulate cognitive overload by assuming patients consider up to $d$ options. Smaller $d$ favors the gradient descent policy because it offers carefully tailored assortments without being too large. (c) While the gradient descent policy maximizes match quality, it achieves worse fairness because it neglects the worst-off patient. In contrast, pairwise and group-based policies achieve higher fairness by only suggesting high-quality matches.
Offering assortments to batches of patients rather than one-shot allows administrators to exercise some control over patient response orders. We apply the batching algorithm from Section 4.1 and analyze the impact of the number of batches, $L$, on the match quality in our real-world scenario. In Figure 8a, we demonstrate that batch offerings allow administrators to trade off logistical burdens for match quality. While offering assortments in two batches can improve match quality by 8% compared to the single-batch setting, the total response time also increases by 74%. Moreover, larger batch sizes lead to diminishing increases in match quality with a near linear increase in the time taken.
5.4.2 Varied Assortment Sizes
Patients might only consider a subset of providers offered due to choice overload (Chernev et al., 2015). To simulate this, we incorporate the maximum assortment size, $d$, to simulate choice overload (see Section 4.2 for details). In Figure 8b, we vary $d\in\{5,10,15,20,25\}$ and find that the gradient descent policy is the best policy across values of $d$, especially for small $d$. This occurs because the gradient descent policy offers smaller assortments compared to the greedy policy, and so restricting assortment size has less of an impact.
5.4.3 Fairness Considerations
To understand how policies trade-off between fairness and match quality, we plot the minimum match quality against the average match quality in Figure 8c (see Section 4.3 for a more thorough discussion of different notions of fairness). In our fairness computation we exclude patients who fail to match, because including them would make the fairness metric always 0 and would not reflect the difference in fairness across policies. We find that the fairest policy, pairwise, also achieves the lowest average match quality. Fairer policies maximize tend to match fewer patients, which keeps the minimum match quality high but lowers the average match quality. If healthcare administrators are focused on fairness as opposed to match quality, then they can select a fairer policy such as the pairwise policy. We compare policies across one definition of fairness for simplicity, but we note that in reality, there are multiple dimensions of fairness that need to be considered (see Section 4.3).
6 Conclusion and Discussion
The patient-provider relationship is key to quality healthcare, yet provider turnover frequently leads to patients being without providers. Our work studies how to re-match patients in such a scenario through assortment optimization. Our study proposes algorithmic approaches for patient-provider matching systems and offers analytical and empirical insights to inform system design. We further offer contributions to one-shot matching under random response order, an understudied phenomenon within the assortment literature. We conclude by providing recommendations for deploying assortment-based systems into practice and discussing extensions of our work.
Recommendations for Deployment
We translate our takeaways into real-world implications:
  1. 1.
    Larger assortments become increasingly useful as patients become more selective because patients are more likely to reject matches, making it important to give more options. Moreover, assortment sizes should be tailored to patient behavior; the choice model and match probability impact which policy performs best.
  2. 2.
    The level of heterogeneity impacts the amount of work needed to design a good algorithm. For example, patients with diverse preferences require less work to find the ideal match, and even relatively simple baselines would work well. However, homogeneous scenarios (e.g., a small city with few providers) require care when matching patients and providers. For example, when patients all prefer the same provider, simple solutions such as the pairwise policy can perform poorly.
  3. 3.
    Healthcare administrators should be aware of the underlying factors that dictate which patients get matched. For example, providers are scarce in rural areas, so assortments might naturally lead to low match rates in rural areas. Algorithm designers should work with administrators during development to minimize these biases.
Model Extensions
We detail several extensions of this model that better capture real-world patient-provider matching. As noted in Section 2.1, we assume that the response time for each patient is i.i.d. A natural extension would be to develop policies that work well independent of the response order distribution. Our experiments also assume that the choice model, $f_{i}$, is constant across patients. However, in reality, patients might have heterogeneous choice models. For example, in the uniform choice model, the value of $p$ could vary across patients. It would be interesting to see how extensions of our policies would perform under heterogeneous choice models.
Extensions to Other Domains
Future work can extend our framework to other scenarios with random response orders. Examples include matching between students and courses, where universities could offer each student an assortment of courses, and between schools and children, where school choice might lead counties to offer each child multiple options (Chen and Sönmez, 2006).
Acknowledgements
We thank Aakash Lahoti, Keegan Harris, and George Chen for their comments. We thank the reviewers at ML4H for their feedback on a preliminary version of this work. We thank Susan Barrett and Karen Hall for their input from a health system perspective. Co-author Raman is supported by an NSF GRFP Fellowship.
Appendix A Extended Experiments
More patients than providers
Figure 9: We compare the match rate across policies when varying the patient/provider ratio, and find that the greedy and gradient descent policies perform best, while the group-based and pairwise policies lag.
We compare the match rate across policies from Example 5 and find that the gradient descent and greedy policies achieve the highest match rate for large patient/provider ratios (see Figure 9). As shown in Section 3.1, the greedy policy achieves the highest match rate for the uniform choice model, so it is unsurprising that it achieves the highest match rate. Beyond that, the high performance of gradient descent matches the trends with match quality in Example 5.
Impact of $p$ and $\theta$
Figure 10: For $p\leq 0.5$, we find that all four policies achieve the same match rate, while for $p\geq 0.75$, we find that gradient descent achieves a slightly lower match rate compared to the greedy, pairwise, and group-based policies.
We extend Example 6 to compare the match rates across policies when varying $p$ and $\theta$ in Figure 10. All non-random policies have similar match rates for $p\leq 0.5$, implying that match quality is better at differentiating policies than match rate. For larger $p$, the gradient descent policy performs slightly worse in match rate, while the greedy, pairwise, and group-based policies achieve the same match rate. As in Example 6, we find that the best-performing policies for match quality also tend to exhibit the highest match rate (in this case, the group-based policy is optimal for both).
Misspecified Choice Models
Figure 11: We evaluate the performance of policies when the observed match probability, $\hat{p}$, differs from the true match probability, $p$. We find that the observed match probability has little impact on the relative ordering of policies, although the group-based policy improves as $p$ and $\hat{p}$ get closer.
Our experiments in Section 5 consider situations with known values of $p$, but real-world scenarios frequently involve misspecified choice models. To model this, we run policies with some observed match probability $\hat{p}$ while letting $p$ be the match probability. This setup allows us to simulate real-world scenarios where match probability is unknown. We compare policies across $\hat{p}\in\{0.1,0.25,0.5,0.75,0.9\}$ and $p\in\{0.1,0.25,0.5,0.75,0.9\}$.
In Figure 11, we demonstrate that most policies perform similarly across $\hat{p}$ for fixed $p$. This trend holds because the random, greedy, and pairwise policies are unimpacted by $\hat{p}$. We find that the group-based policy gets worse as $\hat{p}$ and $p$ differ; for example, when $p=0.9$ and $\hat{p}=0.1$, the group-based policy performs worse than the pairwise policy, while when $\hat{p}=0.9$, the policies perform similarly. We note that the differences are slight even for the group-based policy, so our policies are generally robust to misspecification in match probabilities.
Appendix B Proofs
See 2.3
Proof.
Suppose that $s$ patients are offered the provider $j=1$: $\lVert\mathbf{X}\rVert_{1}=s$. In the uniform choice model, the probability that any of the $s$ patients select the single provider is $p$, so the probability of provider $j=1$ matching is $1-(1-p)^{s}$. By symmetry in the response order, the chance that each patient is selected is equal, and so the match quality is
$(1-(1-p)^{s})\frac{\sum_{i=1}^{N}\theta_{i,1}\mathbf{X}_{i,1}}{s}$ (20)
For fixed $s$, the optimal assortment selects the $s$ largest values of $\theta_{i,1}$. Next, we note that larger values of $s$ increase the match probability $(1-(1-p)^{s})$, but could decrease the average match quality for the selected patient, $\frac{\sum_{i=1}^{N}\theta_{i,1}\mathbf{X}_{i,1}}{s}$. To balance between the two factors, we can iterate through values of $s$ and compute $(1-(1-p)^{s})$ and $\frac{\sum_{i=1}^{N}\theta_{i,1}\mathbf{X}_{i,1}}{s}$. ∎
See 3.1
Proof.
We construct a problem instance where the greedy policy achieves an $\epsilon$ fraction of the optimal match quality. Let $N=M$ and construct $\theta$ as follows: let $\theta_{1,1}=1$, while $\theta_{i,1}=2\Delta$ for $i\neq 1$, where $\Delta\leq\frac{1}{2}$. Let $\theta_{i,j}=\Delta$ for all $i$ and $j\neq 1$.
Let $\pi$ be the policy so that $\pi(\theta)_{i}=\mathbf{e}_{i}$. The expected match quality for this policy is $p(\Delta(N-1)+1)$; for patients 2 to $N$, it achieves an expected match quality of $p(\Delta(N-1))$ in total, while for patient $1$ it achieves an expected match quality of $p$. Because $\pi^{*}(\theta)$ is optimal, we have
$p(\Delta(N-1)+1)=\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]\leq\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]$ (21)
Next, consider the greedy policy where $\pi^{R}(\theta)=\mathbf{1}$. All patients prefer provider $1$, and because of this, all patients have an equal likelihood of matching with provider $1$. Each patient has a $\frac{1}{N}$ chance of matching with provider $1$, and therefore the expected match quality for provider $1$ is $\frac{1}{N}+\frac{N-1}{N}(2\Delta)$. For all the $N-1$ other providers, we receive a reward of $\Delta$ upon matching, and each patient matches with probability $p$; therefore, we receive a total match quality of at most $\frac{1}{N}+2\frac{N-1}{N}\Delta+p(N-1)\Delta$. Taking the ratio of the greedy and optimal policies yields the following:
$\frac{\frac{1}{N}+\frac{2\Delta(N-1)}{N}+p(N-1)\Delta}{p(\Delta(N-1)+1)}$
$\leq\frac{\frac{1}{N}+2\Delta+p(N-1)\Delta}{p(\Delta(N-1)+1)}$
$\leq\frac{\frac{1}{N}+2\Delta+p(N-1)\Delta}{p}$
$\leq\frac{1}{Np}+\frac{2\Delta}{p}+N\Delta$
Finally, letting $N=\frac{3}{\epsilon p}$ and $\Delta=\min(\frac{\epsilon}{3N},\frac{p\epsilon}{6})$ yields that
$\frac{1}{Np}+2\frac{\Delta}{p}+N\Delta\leq\frac{\epsilon}{3}+\frac{\epsilon}{3}+\frac{\epsilon}{3}=\epsilon$ (22)
Therefore, for any choice of $\epsilon$ and $p$, there exists a choice of $\theta$ so the greedy policy is at most an $\epsilon$ approximation. ∎
See 9
Proof.
The pairwise policy constructs an assortment so each patient is only offered one provider, and no provider is offered to more than one patient. Under this scenario, each patient offered a non-zero assortment matches with probability $p$; that is, if $\pi^{P}(\theta)_{i,j}=1$, then the expected match quality is $p\theta_{i,j}$. The match quality is then $\frac{p}{N}\sum_{i=1}^{N}\pi^{p}(\theta)_{i,j}\theta_{i,j}$
Next, we upper bound the performance of the optimal policy. The optimal policy achieves a match quality of
$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]$ (23)
For any order $\sigma$, we have that no provider can be selected by two patients, and each patient can select at most one provider. That is, for any ordering, we can bound the match quality as
$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]\leq\max\limits_{\mathbf{X},\sum_{i=1}^{N}X_{i,j}\leq 1,\sum_{j=1}^{M}X_{i,j}\leq 1}\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}X_{i,j}\theta_{i,j}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}$ (24)
The second inequality comes from the definition of $\pi^{P}(\theta)$. Therefore:
$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]\leq\frac{1}{N}\sum_{i=1}^{N}\pi^{p}(\theta)_{i,j}\theta_{i,j}=p\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{p}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]$ (25)
See 2
Proof.
We will first prove the reverse direction; that if $G$ is a complete digraph, then $\pi$ and $\pi^{P}$ have the same match rate. Consider patient $i$ in a component that is a connected component of size $k$. This means that $\sum_{j}\pi(\theta)_{i,j}=1$. We will prove the set of providers unmatched and offered to patient $i$ is non-empty; that is $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}>0$ for any ordering $\sigma$ and $\mathbf{y}^{(t)}$ with $\sigma_{t}=i$. $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}=k$ when $\mathbf{y}^{(t)}=\mathbf{1}$, and decreases by one each time a provider from $\pi(\theta)_{\sigma_{t}}$ is selected. For any $j$ with $\pi(\theta)_{i,j}=1$, there exists $k-1$ other $i^{\prime}$ with $\pi(\theta)_{i^{\prime},j}=1$ because all patients within the complete graph have provider $j$ on their assortment. Therefore, at most $k-1$ other patients can select providers from $\pi(\theta)_{i}$, and therefore $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}$ can decrease by at most $k-1$ before $\sigma_{t}=i$. Therefore, $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}\geq k-(k-1)=1>0$. Moreover, because each patient has a non-empty assortment, in expectation, $Np$ patients match, and so both $\pi$ and $\pi^{P}$ achieve $Np$ matches.
Next, we will prove the forward direction: if the match rates are equal between $\pi$ and $\pi^{P}$, then $G$ consists of components that are complete digraphs. We note that $\pi^{P}$ achieves a match rate of $Np$ because each patient is offered a non-empty assortment of size one. Therefore, the only way for $\pi$ to achieve the same match rate is to offer each patient a non-empty assortment. To show this, we first consider some node $u$ in the graph, corresponding to a patient. Suppose that patient $i$ has an assortment of size $k$, indicating that there exist $k$ edges from $i$ to some node. Suppose that there exists a node, $i^{\prime}$ such that $(i^{\prime},i)$ is an edge, but $(i,i^{\prime})$ is not. Then consider the ordering that places $i$ after its neighbors and $i^{\prime}$. In such an ordering, suppose $i^{\prime}$ selects $v_{i}$ and each neighbor of $i$, $u$, select $v_{u}$. This results in neither $i$ nor its neighbors being available when $i$ must select a patient, leaving an empty assortment. Therefore, if patient $i$ has $v_{i^{\prime}}$ on its assortment, then $i^{\prime}$ must have $v_{i}$ on its assortment for the policy $\pi$.
Next, we consider the scenario where there exists a node $w$ such that $w$ is a two-hop neighbor of $i$ but not an immediate neighbor of $i$. Suppose $w$ is adjacent to $i^{\prime}$, so that $i^{\prime}$ is on $w$’s assortment. Order the patients such that $w$ comes first, then $i$’s neighbors, then $i$. Let $w$ select $i^{\prime}$, let $i^{\prime}$ select $i$, and let all of $i$’s other neighbors select themselves. This results in an empty assortment for $i$; therefore, when all assortments are nonempty, there must not exist any two-hop neighbors that are not also one-hop neighbors. In other words, every node in $G$ only has one-hop neighbors. Therefore, in any component, all neighbors are connected, and so $G$ consists of components that are complete digraphs. ∎
See 3.1
Proof.
We first compute the expected match quality for the pairwise policy. Each of the $M$ matched patients has a $p$ chance of selecting their assigned provider, and $\theta_{i,j}\leq 1$; therefore, $\mathbb{E}_{\sigma,\theta}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{P}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]\leq pM$.
[…]
$\mathbb{E}[\theta_{i,j}|\text{j is top avail. for i}]=\frac{\sum_{j=1}^{M}\beta_{t}(1-\beta_{t})^{j-1}\mathbb{E}[(\theta_{i})_{(j)}]}{\sum_{j=1}^{M}\beta_{t}(1-\beta_{t})^{j-1}}$ (26)
By Chebyshev’s sum inequality, we have that
$\frac{\sum_{j=1}^{M}\beta_{t}(1-\beta_{t})^{j-1}\mathbb{E}[(\theta_{i})_{(j)}]}{\sum_{j=1}^{M}\beta_{t}(1-\beta_{t})^{j-1}}\geq\frac{1}{M}\sum_{j=1}^{M}\mathbb{E}[(\theta_{i})_{(j)}]=\frac{1}{2}$ (27)
Therefore, summing across all providers gives the ratio as
$\frac{1-(1-p)^{N/M}}{2p}$ (28)
See 2
Proof.
We compute a matrix, $\hat{\mathbf{Y}}$, such that $Y_{i,j}$ corresponds to the probability that patient $i$ is matched with $j$. Then $\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]=\hat{\mathbf{Y}}\cdot\theta$.
We then need to prove that
$(p*g(f(\mathbf{X}))-\hat{\mathbf{Y}})\cdot\theta\leq 0$ (29)
To prove this, first let $\ell_{u_{i,j}}=(p*g(f(\mathbf{X}))_{i,j}-\hat{Y}_{i,j})$.
We will first bound $\sum_{j=1}^{M}\ell_{u_{i,j}}$ for a fixed $i$. Let $\theta_{i,u_{i,1}}\geq\theta_{i,u_{i,2}}\ldots\theta_{i,u_{i,M}}$ for some coefficients $u$. For any $m$, $\sum_{j=1}^{m}\ell_{u_{i,j}}=\sum_{j=1}^{m}p*g(f(\mathbf{X}))_{i,u_{i,j}}-\hat{\mathbf{Y}}_{i,u_{i,j}}$. We can next compute $\sum_{j=1}^{m}p*g(f(\mathbf{X}))_{i,u_{i,j}}$. Recall from Section 3.4 that we can explicitly compute $g(f(\mathbf{X}))_{i,j}$ through the function $h(n)=\frac{1}{N}\sum_{k=1}^{N}(1-p\frac{n-1}{N-1})^{k-1}$. Then by the of definition $g(f(\mathbf{X})_{i,j})$, we have $p\sum_{j=1}^{m}g(f(\mathbf{X}))_{i,j}=p(\sum_{j=1}^{m}X_{i,u_{i,j}}h(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})\prod_{l=1}^{j-1}(1-(\lVert\mathbf{X}_{*,u_{i,l}}\rVert_{1})))$.
Next, we note that $\sum_{j=1}^{m}\hat{\mathbf{Y}}_{i,u_{i,j}}$ represents the probability that any of the top-$m$ options are available and selected. For any provider $j$, the probability that $j$ is available is at least $h(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})$ (see Section 3.4). Therefore,
$\sum_{j=1}^{m}\hat{\mathbf{Y}}_{i,u_{i,j}}$ (30)
[…]
$\geq p\sum_{j=1}^{m}X_{i,u_{i,j}}h(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})\prod_{l=1}^{j-1}(1-(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})$ (32)
Combining our statements for $\hat{\mathbf{Y}}$ and $p*g(f(\mathbf{X}))$ gives that $\sum_{j=1}^{m}X_{i,u_{i,j}}\hat{\mathbf{Y}}_{i,u_{i,j}}\geq\sum_{j=1}^{m}h(\lVert\mathbf{X}_{*,u_{i,m}}\rVert_{1})\prod_{l=1}^{m-1}(1-(\lVert\mathbf{X}_{*,u_{i,m}}\rVert_{1})=\sum_{j=1}^{m}p*g(f(\mathbf{X}))_{i,u_{i,j}}$. Therefore, we have that $\sum_{j=1}^{m}\ell_{u_{i,j}}\leq 0$.
Next, we will show that $(p*g(f(\mathbf{X}))_{i}-\hat{\mathbf{Y}}_{i})\cdot\theta_{i}=\sum_{j=1}^{M}\ell_{u_{i,j}}\theta_{i,j}\leq 0$ for all $i$. We will prove this inductively; first note that $l_{i,u_{i,1}}\theta_{i,u_{i,1}}\leq 0$ Next, assume that $\sum_{j=1}^{M-1}\ell_{u_{i,j}}\theta_{i,u_{i,j}}$, then:
$\sum_{j=1}^{m}\ell_{u_{i,j}}\theta_{i,u_{i,j}}$ (33)
$=\sum_{j=1}^{m-1}\ell_{u_{i,j}}\theta_{i,u_{i,j}}+\ell_{m}\theta_{i,u_{i,m}}$ (34)
$\leq\sum_{j=1}^{m-1}\ell_{u_{i,j}}\theta_{i,u_{i,j}}+\ell_{m}\theta_{i,u_{i,m-1}}$ (35)
$\leq(\sum_{j=1}^{m-1}\ell_{u_{i,j}})(\sum_{j=1}^{m-1}\theta_{i,u_{i,j}}-\theta_{i,u_{i,m-1}})\leq 0$ (36)
Therefore, $p*(g(f(\mathbf{X}))_{i}\cdot\theta_{i})\leq 0$ for all $i$, and so summing across all $i$ gives $p*(g(f(\mathbf{X}))\cdot\theta)\leq 0$
See 2
Proof.
Let $\mathbf{X}=\pi^{P}(\theta)$ for some given $\theta$. Then because $\pi^{P}(\theta)$ is computed by solving a 1-1 bipartite matching problem, we know that $\sum_{i=1}^{N}X_{i,j}\leq 1$ for all $j$ and $\sum_{j=1}^{M}X_{i,j}\leq 1$ for all $i$. Next, we note that $0\leq\lVert\mathbf{X}_{*,j}\rVert\leq 1$, and moreover, if $\mathbf{X}_{i,j}=1$, then $\lVert\mathbf{X}_{*,j}\rVert=1$. Because $h(1)=1$, we have that $f(\mathbf{X})_{i,j}=\mathbf{X}_{i,j}$. Finally, because each row $i$ of $f(\mathbf{X})_{i,j}$ has at most one $j$ with $\mathbf{X}_{i,j}=1$, and because $\bar{\mathbf{X}}_{i,u_{k}}=f(\mathbf{X})_{i,u_{k}}p\prod_{k^{\prime}=1}^{k-1}(1-f(\mathbf{X})_{i,u_{k}})$, we have that $g(f(\mathbf{X}))=f(\mathbf{X})=p\mathbf{X}$. Therefore, $g(f(\mathbf{X}))\cdot\theta$ is exactly p times the value of the bipartite match. Finally,
$\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]=p\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}=p(f(g(\mathbf{X}))\cdot\theta).$ (37)
Therefore, the expected match quality for the pairwise policy $\pi^{P}(\theta)$ is the same as $p(\mathbf{X}\cdot\theta)$. ∎
See 4.1
Proof.
We first note that $G$ is acyclic; if $\theta_{i,v_{i}}\leq\theta_{i,v_{i^{\prime}}}$ and $\theta_{i^{\prime},v_{i^{\prime}}}\leq\theta_{i^{\prime},v_{i}}$, then swapping $v_{i}$ and $v_{i^{\prime}}$ improves the pairwise policy $\pi^{P}(\theta)$. We note that $\pi^{P}(\theta)$ is the optimal solution for $\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)\theta_{i,j}$ by definition, and so such a swap would violate this property, implying that $G$ is acyclic.
Next, suppose that $\sigma^{*}$ is not a reverse topological ordering. Then there exists nodes $i$ and $i^{\prime}$, so that $i$ comes before $i^{\prime}$ in $\sigma^{*}$ and there exists an edge from $i$ to $i^{\prime}$. Consider the following cases:
  1. 1.
    Case 1: There exists no node $w$, between $i^{\prime}$ and $i$ in $\sigma$ so that $w$ and $i^{\prime}$ share an edge (in either direction). Then $i$ and $i^{\prime}$ can be placed in sequence ($i$ comes immideatly before $i^{\prime}$) without impacting the expected reward. This is because all nodes $w$ between $i$ and $i^{\prime}$ have no impact on the preferences of $i^{\prime}$ because $\theta_{i^{\prime},v_{i^{\prime}}}\geq\theta_{i^{\prime},v_{w}}$. Moreover, swapping $i$ and $i^{\prime}$ can only increase the expected match quality; placing $i$ before or after $i^{\prime}$ has no impact on the match quality for patient $i^{\prime}$ (as $\theta_{i^{\prime},v_{i^{\prime}}}\geq\theta_{i^{\prime},v_{i}}$), but placing $i$ after $i^{\prime}$ can increase the expected match quality if $v_{i}$ goes unmatched.
  2. 2.
    Case 2: Suppose there exists a node $w$, such that there is an edge from $w$ to $i^{\prime}$ Then the pair $(w,i^{\prime})$ violates the reverse topological order, and so we can recurse on this. Note that this is a smaller length pair of nodes within the ordering $\sigma^{*}$. Similarly if there is an edge from $i$ to $w$, then $(i,w)$ violates the reverse topological order ,and so we can recurse on this
  3. 3.
    Case 3: Suppose that there exists a node $w$ such that there are edges from $i^{\prime}$ to $w$ and $w$ to $i$. Then this creates a cycle, breaking the acyclic property mentioned earlier.
  4. 4.
    Case 4: Suppose that there exists a node $w$ such that there is an edge from $i^{\prime}$ to $w$. There exists no path from $i^{\prime}$ to $i$ due to the acyclic nature of the graph. Additionally, $w$ does not have an edge to $i$ or any of the ancestors of $i$. Move $i^{\prime}$ and all of its descendants before $i$ in the graph; there are no edges from the descendants of $i^{\prime}$ to $i$. Therefore, this flips the order of $i$ and $i^{\prime}$ without impacting any of the descendants; because only $i$ and $i^{\prime}$ are impacted by this change (as the descendants of $i^{\prime}$ are not impacted by $i$), such a change can only increase the expected match quality, as $i$ can potentially match with $v_{i^{\prime}}$.
  5. 5.
    Case 5: Suppose there exists a node $w$ such that there exists an edge from $w$ to $i$. Due to the acyclic nature of the graph, there are no edges from $i^{\prime}$ (or any of its ancestors) to $w$ or any of the ancestors of $i$. Now consider moving $i$ and its ancestors after $i^{\prime}$. Because there are no edges from $i^{\prime}$ to the ancestors of $i$, the result of the assortment from the ancestors of $i$ has no impact on $i^{\prime}$. Therefore, while patient $i$ can now potentially match with provider $v_{i^{\prime}}$, none of the ancestors of $i$ are negatively impacted, improving aggregate match quality.
Our cases cover every scenario for nodes between $i$ and $i^{\prime}$, and in all cases, we can make a minor change to the order $\sigma^{*}$ to bring it closer to a reverse topological ordering while improving match quality or recurse upon a subset of the ordering. ∎
See 4.1
Proof.
Let $\sigma^{*}$ be the optimal ordering and let $\sigma$ be any ordering from batches $b_{1},b_{2},\ldots,b_{L}$. Then we will prove that for any node $i$, the order of descendants of $i$ will be the same in $\sigma^{*}$ and $\sigma$.
To prove this, consider a node $i$. Let the descendants of node $i$ in the optimal ordering be $d_{1},d_{2},\ldots,d_{i}$. We note that in $\sigma$, each $d_{i^{\prime}}$ is in a separate partition by properties of $b_{k}$. Additionally, because $b_{1},b_{2},\ldots,b_{L}$ is a partition of $\sigma^{*}$, the order of $d_{1},d_{2},\ldots,d_{i}$ is maintained. Therefore, for any node, the same set of ancestors is maintained through the partition $b_{k}$.
Next, we will show that the descendants of patient $i$ dictates what patient $i$ selects in $\pi^{A}$. Let $Z_{1},Z_{2},\ldots,Z_{N}$ be a set of Bernoulli random variables, each of which is 1 with probability $p$. If $Z_{t}=1$ then patient $\sigma_{t}$ will select the highest available provider in their assortment, and if $Z_{t}=0$, then patient $\sigma_{t}$ will select no provider. Additionally, let $V_{1},V_{2},\ldots,V_{N}$ be a set of random variables, where $Z_{t}$ denotes the provider selected by patient $\sigma_{t}$ ($V_{t}=0$ if $Z_{t}=0$). Note that $V_{1}$ is only a function of $Z_{1}$; if $Z_{1}=0$, then $V_{1}=0$, and otherwise, $V_{1}=\operatorname*{arg\,max}_{j}\theta_{\sigma_{1},j}$. Next, let $D_{\sigma_{t}}\subseteq[N]$ be the set of descendants for patient $\sigma_{t}$; that is, $d\in D_{\sigma_{t}}$ means there exists a path from $\sigma_{t}$ to $d$ in $G$. Suppose that $V_{\sigma_{t^{\prime}}}=f(\{Z_{d}\}_{d\in D_{\sigma_{t^{\prime}}}})$ for all $t^{\prime}\leq t$. Then patient $\sigma_{t+1}$ will only select some $j$ so $\theta_{\sigma_{t+1},j}\geq\theta_{\sigma_{t+1},v_{\sigma_{t+1}}}$. By the construction of the graph $G$, this corresponds to edges connected to patient $\sigma_{t+1}$, and so $V_{t+1}=g(\{V_{d}\}_{d\in D_{\sigma_{t+1}})}$. Next, note that $V_{d}=f(\{Z_{d^{\prime}}\}_{d^{\prime}\in D_{d}})$. Therefore, we can rewrite $V_{t+1}$ as
$V_{t+1}=g(\{f(\{Z_{d^{\prime}}\}_{d^{\prime}\in D_{d}})\}_{d\in D_{t+1})}.$ (38)
We finally collect like terms and note that the set of descendants of the children of $\sigma_{t+1}$ is the set of descendants of $\sigma_{t+1}$. Therefore, we rewrite $V_{t+1}$ as
$V_{t+1}=g(\{Z_{d}\}_{d\in D_{t+1}})$ (39)
Therefore, $V_{t+1}$ only depends on the descendants of $t+1$. Therefore, if the ordering of descendants is fixed, and $Z_{1},Z_{2},\ldots,Z_{N}$ are decided a priori, then $V_{\sigma_{t}}$ is also fixed. The ordering of descendants is the same between $\sigma\sim S(b_{1},b_{2},\ldots,b_{L})$ and $\sigma^{*}$, so both achieve the same match quality. ∎